### v2

# Constraining Super-Heavy Dark Matter with the KM3-230213A Neutrino Event

Abstract
Recently, the KM3NeT collaboration detected an astrophysical neutrino event, KM3-230213A, with an energy of approximately $220\penalty 10000\ \rm PeV$, providing unprecedented insights into the ultra-high-energy Universe. In this study, we introduce a novel likelihood framework designed to leverage this event to constrain the properties of super-heavy dark matter (SHDM) decay. Our approach systematically integrates multi-messenger constraints from galactic and extragalactic neutrino flux measurements by IceCube, the absence of comparable neutrino events at IceCube and Auger observatories, and the latest gamma-ray experiment upper limits. Our findings impose the most stringent constraints to date, placing a lower bound on the SHDM lifetime at $\gtrsim 5\cdot 10^{29}-10^{30}\rm s$. Importantly, we identify, for the first time, the significant potential of galactic neutrino flux measurements in advancing dark matter research. Future investigations targeting astrophysical neutrinos originating from the Galactic Center at energies above $10\penalty 10000\ \rm PeV$ will be crucial, not only for understanding the origin of the cosmic-ray knee but also for exploring the possible contributions of super-heavy dark matter to our Universe.
I Introduction
Dark matter (DM) is a cornerstone of the standard cosmological model Aghanim and others (2020), yet its fundamental nature remains elusive. Over the years, extensive efforts have been devoted to the search for DM particle candidates, spanning a wide mass range from ultra-light ($\sim 10^{-5}\,\rm eV$) to super-heavy ($\sim 10^{21}\,\rm eV$) scenarios Bertone and Tait (2018); Kahlhoefer (2017); Billard and others (2022); Argüelles et al. (2023); Maity et al. (2021); Guépin et al. (2021). In this context, Multi-Messenger Astroparticle Physics provides unique opportunities to probe the properties of DM, as DM annihilation or decay should leave observable signatures in gamma-ray, neutrino, and cosmic-ray data Ambrosone et al. (2023); Aiello and others (2025a); Ajello and others (2015); Abbasi and others (2025b); Abdul Halim and others (2024); Abreu and others (2023b, a). On the night of February $13^{\mathrm{th}}$, 2023, the KM3NeT/ARCA Cherenkov neutrino telescope in the Mediterranean Sea detected a cosmic neutrino with an energy of approximately $220\,\rm PeV$, designated as the KM3-230213A event Aiello and others (2025b). This detection may represent a new component in the high-energy neutrino sky, not previously observed by either the IceCube or Pierre Auger Observatories Abbasi and others (2025a); Aab and others (2019); Abdul Halim and others (2025). It has been speculated that the KM3NeT-230213A event may provide a hint of physics beyond the Standard Model (BSM) Zantedeschi and Visinelli (2025); Adriani and others (2025b); Barman et al. (2025); Dvali et al. (2025b); Jho et al. (2025); Klipfel and Kaiser (2025); Telalovic and Bustamante (2025); Jiang and Huang (2025); Alves et al. (2025); Turiaci and Wu (2025); Narita and Yin (2025); Kohri et al. (2025); Brdar and Chattopadhyay (2025); Boccia and Iocco (2025); Dvali et al. (2025a); Amelino-Camelia et al. (2025); Borah et al. (2025); Khan et al. (2025b); Murase et al. (2025); Bertólez-Martínez et al. (2025); Airoldi et al. (2025); Anchordoqui et al. (2025); Dev et al. (2025); Farzan and Hostert (2025); Baker et al. (2025); Khan et al. (2025a); Choi et al. (2025); Sakharov et al. (2025), possibly associated with phenomena such as primordial black holes, quantum gravity effects, or the decay of dark matter. In particular, DM candidates with masses above $\sim 10^{8}\,\rm GeV$, collectively referred to as super-heavy dark matter (SHDM), have been proposed as a natural solution to several persisting cosmological problems (see, e.g., Baer et al. (2015); Bertone et al. (2005) for reviews). However, these interpretations—particularly the SHDM decay scenario—do not account for the fact that the expected signal from SHDM decay should peak toward the Galactic Center, whereas the coordinates of KM3-230213A ($l_{\rm event}=216.1^{\circ}$, $b_{\rm event}=-11^{\circ}$) lie well away from it Aiello and others (2025b). Furthermore, given the energy of the event, cosmogenic neutrinos remain the most viable explanation Aloisio et al. (2015); Adriani and others (2025c); Abdul Halim and others (2023); Cermenati et al. (2025). In this work, assuming that the KM3-230213A event originates from a diffuse astrophysical background, we place limits on the SHDM hypothesis. Unlike previous studies Barman et al. (2025); Dvali et al. (2025b); Jho et al. (2025); Jiang and Huang (2025); Narita and Yin (2025); Kohri et al. (2025); Brdar and Chattopadhyay (2025); Borah et al. (2025), we develop a comprehensive likelihood framework that incorporates constraints from ultra-high-energy (UHE) gamma-rays Abdul Halim and others (2025); Apel and others (2017); Chantell and others (1997); Fomin et al. (2017); Abbasi and others (2019, 2021) as well as neutrino fluxes and upper limits reported by the IceCube collaboration Abbasi and others (2023, 2022). Moreover, we consistently model both the Galactic contribution, adopting a representative dark matter halo profile, as well as the uniform extragalactic contribution, ensuring a robust treatment of the expected signal across all relevant spatial components. We place the strongest constraints to date on the DM lifetime $(\gtrsim 5\times 10^{29}-10^{30}\,\rm s)$. This outcome highlights the critical importance of a global approach for unveiling the properties of a specific dark matter model. Looking ahead, the extension of the galactic neutrino spectrum up to $\sim 10$$100\,\rm PeV$ will be invaluable in constraining such scenarios, as the Galactic DM-induced contribution is expected to be predominant along the Galactic Plane.
Figure 1: $95\%$ CL lower limit of the SHDM in terms of $M_{\chi}$ in the mass range $10^{7}-10^{17}\,\rm GeV$. The orange band represent the excluded regions obtained by our analysis. Left: It refers to $\chi\longrightarrow\nu\bar{\nu}$ and it is compared to the limits obtained by Chianese et al. (2021). Right: It refers to $\chi\longrightarrow b\bar{b}$ and it is compared with the limits obtained by Das et al. (2023); Cao and others (2022)
II SHDM Decay Gamma-Ray and Neutrino Fluxes
The total flux arising from SHDM decay is given by the sum of Galactic and extragalactic contributions Chianese et al. (2019).
The Galactic, per solid-angle flux is given by
$\begin{split}\frac{d\phi_{G,i}}{dEd\Omega}\left(E,M_{\chi},\tau_{\chi}\right)=\frac{1}{4\pi M_{\chi}\tau_{\chi}}\frac{dN_{i}}{dE}(E)\\ \times\int ds\,\rho(s,l,b)\,e^{-s/\lambda_{i}(l,b)}\penalty 10000\ ,\end{split}$ (1)
where $i=\gamma,\nu$ denotes either gamma rays or neutrinos, $M_{\chi}$ and $\tau_{\chi}$ are respectively the mass and lifetime of the DM particle $(\chi)$. $d\Omega=\cos b\,db\,dl$ is the solid angle element, $\rho(s,l,b)$ is the dark matter energy density (in units of $\rm GeV\,cm^{-3}$), depending on the distance $s$ along the line of sight and Galactic coordinates $(l,b)$. The term ${dN_{i}(E)}/{dE}$ is the gamma-ray or neutrino energy spectrum per DM decay, computed using the public code HDMSpectra Bauer et al. (2021), which accounts for electroweak corrections Ciafaloni et al. (2011). $\lambda_{i}(l,b)$ is the effective attenuation length of gamma rays or neutrinos. Since neutrinos propagate essentially unimpeded up to very high redshift ($z\gtrsim 100$), we set $\lambda^{\nu}\rightarrow+\infty$. For gamma rays, the main attenuation is due to pair production on the cosmic microwave background (CMB) Das et al. (2023). For the Galactic DM energy density, we adopt the Navarro-Frenk-White (NFW) profile Navarro et al. (1996), exploiting the parameter values reported in Chianese et al. (2019). The Galactic DM density profile remains uncertain and can alternatively be described by other parametrizations, such as the Burkert profile Burkert (1995); Salucci and Burkert (2000). However, we have verified that this uncertainty has a negligible effect on our results ($\lesssim 5\%$), and we thus neglect it in the following.
The extragalactic DM distribution also contributes to the overall flux. The differential flux per solid angle from extragalactic decays reads Chianese et al. (2019)
$\frac{d^{2}\phi_{EG,i}}{dE\,d\Omega}(E)=\frac{\Omega_{\chi}\rho_{\rm cr}\,c}{4\pi M_{\chi}\tau_{\chi}H_{0}}\int_{0}^{z_{\rm max}}\frac{dz}{E(z)}\,\frac{dN_{i}}{dE}\big(E(1+z)\big),$ (2)
[…]
III On the Direction of the KM3-230213A Event
In this section, we demonstrate that the direction of KM3-230213A strongly disfavors a galactic origin, specially in the context of SHDM decay, by exploiting the concept of fraction visibility $F$ of neutrino telescopes. The visibility of an experiment is defined as the fraction of the day a source, in a given equatorial declination, lies within a given interval of zenith angle Aiello and others (2024a). This limits the field of view of the experiment and it is simply dependent on the location of the detector. At ultra high energies $(E\geq 1-10\,\rm PeV)$, neutrinos can be observed specifically through down-going and horizontal events because of the Earth absorption Aiello and others (2025b); Abbasi and others (2025a). For KM3NeT/ARCA and IceCube detectors, the corresponding zenith angle range is $0^{\circ}\leq\theta\leq 100^{\circ}$ Aiello and others (2024b). For the Pierre Auger Observatory, we consider $60^{\circ}\leq\theta\leq 95^{\circ}$ according to the range used in Aab and others (2019). Fig. 2 reports the fractional visibility as a function of the declination for the experiments, evaluated through the public code available at https://zenodo.org/records/10955399 KM3NeT and van Eeden (2024); Aiello and others (2024a). Remarkably, the Galactic Center is $40\%$ of the time observable in the down-going/horizontal sky for KM3NeT/ARCA. Consequently, The Galactic Center should anyway dominate in case of SHDM decay, where the flux is principally concentrated. To quantify this effect further, Figs. 3 and 4 show 2D maps in galactic coordinates of the product between the visibility and the line-of-sight integral $\int ds,\rho(s,l,b)$, representing the expected contribution from Galactic dark matter decay for KM3NeT/ARCA, IceCube, and the Pierre Auger Observatory, respectively.
Figure 2: The fractional visibility $F$ as a function of the equatorial declination for KM3NeT/ARCA (blue line), IceCube (orange line) and Pierre Auger Observatory (green line). The vertical dashed red line refers to the Galactic center declination.
Figure 3: 2D map in galactic coordinates of the visibility $\rm F$ $\times$ the galactic DM distribution flux for KM3NeT/ARCA, considering $0^{\circ}\leq\theta\leq 100^{\circ}$. The blank region corresponds to the blind spots, while the golden and red stars respectively are the Galactic center and the location of KM3-230213A event.
Figure 4: The same as Fig. 3. Left The IceCube case, considering $0^{\circ}\leq\theta\leq 100^{\circ}$. Right: The Pierre Auger case, considering $60^{\circ}\leq\theta\leq 95^{\circ}$.
In all cases, the maps are clearly dominated by the Galactic Center. Although detector response and event selection are not included, this trend highlights the central role of the Galactic Center in shaping the expected signal.
IV Constraints and Analysis Setup
Our analysis evaluates the constraining power of gamma-rays and neutrinos fluxes into constraining the SHDM decay scenario, by enforcing the following constraints:
  • The total neutrino flux (Galactic plus extragalactic) must not exceed the diffuse astrophysical flux measured by the IceCube Collaboration Abbasi and others (2022).
  • The Galactic neutrino component should be bounded by the flux reported by IceCube’s dedicated analysis of the Galactic plane Abbasi and others (2023).
  • The SHDM decay lifetime must be compatible with non-observation of a steady neutrino flux by IceCube Abbasi and others (2025a) and the Pierre Auger Observatory Aab and others (2019) in the energy range relevant for KM3-230213A.
  • The associated gamma-ray flux must not overshoot the integrated upper limits set by gamma-ray and cosmic ray experiments Abdul Halim and others (2025); Apel and others (2017); Chantell and others (1997); Fomin et al. (2017); Abbasi and others (2019, 2021).
To consistently incorporate these constraints, we define the following likelihood function:
$\mathcal{L}(M_{\chi},\tau_{\chi})=\$$\mathcal{P}\left[\sum_{i}N_{i}(M_{\chi},\tau_{\chi})+N_{\rm back}^{\rm astro},1\right]$
$\times\exp\left[-\frac{\chi^{2}_{\gamma}(M_{\chi},\tau_{\chi})}{2}\right]$
$\times\exp\left[-\frac{f^{2}_{\rm G}(M_{\chi},\tau_{\chi})}{2\sigma_{\rm G}^{2}}\right]$
$\times\exp\left[-\frac{\chi^{2}_{\rm D}(M_{\chi},\tau_{\chi})}{2\sigma_{\rm D}^{2}}\right]$ (3)
where the sum is over
$i=\{\text{KM3NeT/ARCA, IceCube, Pierre Auger}\}\penalty 10000\ .$
The function $\mathcal{P}(\sum_{i}N_{i},1)$ denotes the Poisson probability for the combined detectors to observe one event given the expected sum $\sum_{i}N_{i}(M_{\chi},\tau_{\chi})$. This approach effectively treats the three experiments as a single, composite detector, with the recorded event in KM3NeT/ARCA representing a detection in one segment of the overall exposure. A similar statistical treatment was adopted in Adriani and others (2025d) to determine the isotropic flux corresponding to the detection of KM3-230213A. To compute the expected number of events in the energy interval above $10\,\rm PeV$ (denoted as $\Delta E$), we include both extragalactic and Galactic DM contributions. For the extragalactic flux, the expected number of events for each experiment is given by:
$N^{\rm EG}_{i}=4\pi T_{i}\int_{\Delta E}A^{\rm eff}_{i}(E)\frac{d^{2}\phi^{\nu}_{\rm EG}}{dE\,d\Omega}(E;M_{\chi},\tau_{\chi})dE\,$, (4)
where $T_{i}$ and $A^{\rm eff}_{i}(E)$ denote, respectively, the experimental operational time and the energy–dependent effective area of each experiment. For KM3NeT/ARCA we adopt $T_{i}=335\,\mathrm{days}$, while for IceCube we use $T_{i}=12.6\,\mathrm{yr}$, with effective areas taken from Aiello and others (2025b); Abbasi and others (2025a). For the Pierre Auger Observatory, we employ the exposure reported in Aab and others (2019). For the Galactic component, we use
$N^{\rm G}_{i}=T_{i}\int_{\Delta E}A_{i}^{\rm eff}(E)\frac{d\phi^{\nu}_{\rm G}}{dEd\Omega}(E;M_{\chi},\tau_{\chi},\Omega)F_{i}(\Omega)dEd\Omega\penalty 10000\ ,$ (5)
where $F_{i}(\Omega)$ is the fractional visibility for each experiment, accounting for a limited field of view (FoV) Argüelles et al. (2023). The last three terms in Eq. (IV) are Gaussian penalty factors, each enforcing the corresponding observational constraint described above. The gamma-ray term, $\chi_{\gamma}^{2}(M_{\chi},\tau_{\chi})$, follows the approach of Ref. Chianese et al. (2021) and is defined as
$\chi_{\gamma}^{2}(M_{\chi},\tau_{\chi})=\sum_{j}\left(\frac{\Phi^{\rm G}_{\gamma}(>E_{j};M_{\chi},\tau_{\chi})}{\sigma_{j}}\right)^{2},$ (6)
where the sum runs over all 95% CL upper limits from the various gamma-ray experiments, reported in Abdul Halim and others (2025); Apel and others (2017); Chantell and others (1997); Fomin et al. (2017); Abbasi and others (2019, 2021). Each uncertainty is taken as half the reported upper limit, i.e., $\sigma_{j}=0.5\Phi_{95\%\rm UL}$. Here, $\Phi_{\gamma}(>E_{j})$ is the integrated, angle-averaged gamma-ray flux above energy $E_{j}$, as gamma-ray experiments constrain the energy-integrated flux per solid angle. For the diffuse neutrino flux constraint, we account for the fact that IceCube has not observed any events in the 51005-100 PeV energy range and has set a 90% CL upper limit of $\Phi^{\rm IC}_{90\%}\leq 7.79\times 10^{-16}\rm cm^{-2}s^{-1}sr^{-1}$ Abbasi and others (2022). The corresponding penalty term is
$\chi^{2}_{D}=\left(\frac{\Phi_{\nu}^{\rm G+EG}(>E_{\nu};M_{\chi},\tau_{\chi})}{\sigma_{D}}\right)^{2},$ (7)
where $\sigma_{D}=0.5\Phi^{\rm IC}_{90\%}$, and the total expected neutrino flux is obtained by integrating between 5 PeV and 100 PeV the sum of the Galactic (Eq. 1) and Extra-Galactic (Eq. 2) contribution to the angle-averaged diffuse flux. IceCube has also provided a measurement of the Galactic neutrino flux using a template likelihood analysis of cascade (shower) events in the energy range $\sim 1-10^{3}$ TeV Abbasi and others (2023). The measured neutrino intensity at $E=1,\rm PeV$ is $E^{2}\phi^{\rm meas}_{\nu+\bar{\nu}}(E)\sim 3.89\times 10^{-9}\rm GeVcm^{-2}s^{-1}$, integrated over the sky, with a relative statistical uncertainty reported of $\sigma_{G}\simeq 30$% Abbasi and others (2023). To quantify the corresponding Gaussian penalty, we define the Galactic fractional DM contribution to the measured flux at 1 PeV and require that this fraction $(f_{G})$ be consistent with zero within the stated uncertainty.
$N_{\rm back}^{\rm astro}$ represents the contribution from a potential astrophysical background to the total number of neutrino events observable by the detectors. This contribution is both theoretically and observationally well-motivated, though it is subject to substantial uncertainties that could significantly influence our analysis Ehlert et al. (2024); Aloisio et al. (2015); Aiello and others (2025b); Condorelli et al. (2023); Rossoni et al. (2025); Winter and Lunardini (2023); Alves Batista et al. (2019); Heinze et al. (2019); Adriani and others (2025a). In order to account for it while remaining both conservative and model–independent, we adopt $N_{\rm back}^{\rm astro}=1$. This choice corresponds to attributing the KM3-230213A event to the astrophysical background. This assumption is conservative for several reasons: higher values of $N_{\rm back}^{\rm astro}$ would result in even stronger constraints, and astrophysical backgrounds—such as cosmogenic or in-source neutrinos—could also contribute to the Gaussian penalty terms in the likelihood, further tightening the limits. Future analyses might be able to rely on realistic astrophysical predictions, provided that observational evidence for an astrophysical flux at these energies becomes available. We define a Test Statistic $(\rm TS)$ as
$\mathrm{TS}(M_{\chi},\tau_{\chi})=2\ln\left(\frac{\mathcal{L}(\bar{\tau}_{\chi}\longrightarrow+\infty)}{\mathcal{L}(M_{\chi},\tau_{\chi})}\right)\penalty 10000\ ,$ (8)
and derive a one-sided 95% confidence level (CL) lower limit on the dark matter lifetime as a function of its mass by imposing ${\rm TS}=2.71$ Cao and others (2022); Cowan et al. (2011).
V Results
Figure 1 shows the resulting lower limits on $\tau_{\chi}$ as a function of $M_{\chi}$. The left and right panels refer to $\chi\rightarrow\nu\bar{\nu}$ and $\chi\rightarrow b\bar{b}$ channels, respectively. In general, the decay channel $\chi\rightarrow b\bar{b}$ is more strongly constrained due to its higher gamma-ray yield. Remarkably, even under our conservative assumptions, the KM3-230213A event allows us to set the most stringent constraints to date on $\tau_{\chi}$ for both decay channels considered. Specifically, in the case of $\chi\longrightarrow b\bar{b}$, we find an improvement of up to a factor of $\sim 1.5$ for $M_{\chi}\lesssim 10^{11},\rm GeV$ compared to Ref. Das et al. (2023), while our results are comparable at higher masses. Furthermore, for $M_{\chi}\lesssim 10^{9}\,\rm GeV$, our constraints surpass those obtained by the LHAASO collaboration Cao and others (2022). Our findings reinforce previous results Abreu and others (2023b, a); Bérat et al. (2022); Das et al. (2023), underscoring the power of gamma-ray observations in constraining a broad class of beyond-the-Standard-Model scenarios. Further improvements in gamma-ray upper limits at the highest energies Castellina (2019) will provide complementary information and play a decisive role in probing or excluding SHDM-induced signatures even further. In the case of $\chi\rightarrow\nu\bar{\nu}$, the limits improve by up to a factor of $\sim 3$ compared to Ref. Chianese et al. (2021), and by an even larger factor relative to those reported in Ref. Argüelles et al. (2023). This enhanced sensitivity arises from a higher neutrino yield associated with this decay channel. The galactic neutrino flux, in the energy range $\gtrsim 10\,\rm PeV$, is crucial to falsify the SHDM decay hypothesis even further. Therefore, future experiments such as GRAND, RNG-0, IceCube Gen 2 and POEMMA promise unprecedented sensitivity, potentially enabling the extension of these constraints to higher energies Álvarez-Muñiz and others (2020); Aartsen and others (2021); Olinto and others (2021); Agarwal et al. (2025).
VI Conclusions
In this manuscript, we have utilized the highest-energy neutrino ever detected, KM3-230213A, to establish the most stringent constraints on the SHDM decay scenario to date. Even adopting a conservative approach, we derive robust lower bounds on the dark matter lifetime, $\tau_{\chi}\gtrsim 5\cdot 10^{29}-10^{30}\rm s$ for $10^{7}\,\rm GeV\lesssim m_{\chi}\lesssim 10^{17}\,\rm GeV$, slightly varying based on the decay channel considered. Future analyses employing detailed spatial distribution templates are expected to further refine these constraints, thereby placing neutrino astronomy at the forefront of probing SHDM models.
We anticipate that imminent advancements in neutrino observatories – most notably KM3NeT/ARCA and IceCube-Gen2 – with their enhanced exposure and sensitivity toward the Galactic neutrino background, complemented with upcoming ultra-high-energy gamma-ray observations, such as those from AugerPrime Castellina (2019), will jointly play a crucial role in refining constraints on the SHDM decay hypothesis, potentially providing decisive evidence against hadronic decay channels or tightening constraints on leptonic channels.
Acknowledgements.
Acknowledgments: AA acknowledges the support of the project “NUSES - A pathfinder for studying astrophysical neutrinos and electromagnetic signals of seismic origin from space” (Cod. id. Ugov: NUSES; CUP: D12F19000040003). The work of RA and CE has been partially funded by the European Union – Next Generation EU, through PRIN-MUR 2022TJW4EJ and by the European Union – NextGenerationEU under the MUR National Innovation Ecosystem grant ECS00000041 – VITALITY/ASTRA – CUP D13C21000430001.